OpenStudy (anonymous):

Prove the identity: secx+cosx/secx-cosx = 1+cos^2x/sin^2x

OpenStudy (lυἶცἶ0210):

Have you tried converting them into sin and cos?

OpenStudy (lυἶცἶ0210):

\(\LARGE \frac{1}{cosx}+\frac{cosx}{1}*\frac{cosx}{1} -cosx =1+\frac{cos^2x}{sin^2x} \)

OpenStudy (anonymous):

Thank you for your help apparently the program we have to use for school did not like the way I was simplifying the fraction. I just wanted to make sure I was doing it correctly.